calculate the change in melting point of ice when it is subjected to a pressure of 100 atm. Density of ice is 0.917g/c.c and latent heat of ice is 336J/g
Answer :72.26k
1
Expert's answer
2019-04-08T09:43:47-0400

Clausius–Clapeyron relation:

"\frac{dP}{dT}=\frac{Q}{T \Delta V}"

where"Q=336\, J/g" - latent heat of ice, "\Delta V = \left(\frac{1}{0.917}-1\right) \, cm^3 /g" is the specific volume change of unit mass of the phase transition.As we know, melting point of ice at pressure 1atm at temperature

"T_0=0^\circ C = 273.15K"


Assume no (or tiny) density of ice and water dependence on temperature and pressure.

Therefore,


"\int dP=\int \frac{QdT}{T \Delta V} \Rightarrow \frac{T}{T_0}=e^{-\frac{\Delta P \Delta V}{Q}} = e^{-\frac{99\,atm \, 0.0905 \, cm^3 /g}{336\,J/g}}=0.997337"

Hence

"\Delta T=T-T_0 = 272.423 - 273.15=-0.72749 K"